# Give Me FP32 or Give Me Death? Challenges and Solutions for Reproducible Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) are now integral across various domains and have demonstrated impressive performance. Progress, however, rests on the premise that benchmark scores are both accurate and reproducible. We demonstrate that the reproducibility of LLM performance is fragile: changing system configuration such as evaluation batch size, GPU count, and GPU version can introduce significant difference in the generated responses. This issue is especially pronounced in reasoning models, where minor rounding differences in early tokens can cascade into divergent chains of thought, ultimately affecting accuracy. For instance, under bfloat16 precision with greedy decoding, a reasoning model like DeepSeek-R1-Distill-Qwen-7B can exhibit up to 9% variation in accuracy and 9,000 tokens difference in response length due to differences in GPU count, type, and evaluation batch size. We trace the root cause of this variability to the non-associative nature of floating-point arithmetic under limited numerical precision. This work presents the first systematic investigation into how numerical precision affects reproducibility in LLM inference. Through carefully controlled experiments across various hardware, software, and precision settings, we quantify when and how model outputs diverge. Our analysis reveals that floating-point precision—while critical for reproducibility—is often neglected in evaluation practices. Inspired by this, we develop a lightweight inference pipeline, dubbed LayerCast, that stores weights in 16-bit precision but performs all computations in FP32, balancing memory efficiency with numerical stability. Code is available at https://github.com/nanomaoli/llm_reproducibility.
@@ 1 Introduction
We highlight the significance of this issue for several critical reasons. First, many researchers report benchmark performance based on a single inference run with a fixed random seed to reduce computational cost. However, as shown in Figure 1. This practice can lead to misleading conclusions about model performance. Second, even when performing random sampling with multiple independent runs, the averaged results can still vary significantly due to hardware and system-level non-determinism. Moreover, when researchers report standard deviations without accounting for this numerical non-determinism, they risk severely overestimating a model’s true uncertainty, since the reported variance reflects a mixture of intrinsic model uncertainty and variance introduced by finite numerical precision. When results cannot be exactly reproduced, it becomes difficult to distinguish whether improvements are from better methods or merely random variation.
@@
  • An optimized FP32 inference pipeline. We propose an optimized inference pipeline called LayerCast, which performs all computations in FP32 while retaining model weights in BF16 precision. This approach effectively balances memory efficiency and reproducibility. We release it as a patch to vLLM and and can be used with just a few lines of code change.
2 Preliminary
@@ 2.1 Current Practices on LLM Reproducibility
There are two widely adopted evaluation strategies. The first one is greedy decoding, where temperature is set to zero and model always select the token with highest probability. The second strategy employs random sampling with a non-zero temperature, and evaluates performance using the Pass@K metric chen2021evaluating . Below we introduce the commonly adopted experimental setting for the reproducibility.
@@
Deterministic libraries and Random Seeds. Random seeds in LLM generation control the pseudo-random number generator that selects tokens when using non-greedy decoding. It is a standard practice to fix random seeds; however, this is not always sufficient for ensuring full determinism in greedy decoding. Even with a fixed seed, differences in computation order can alter the sequence of the pseudo-random number generator hochlehnert2025sober . Meanwhile, frameworks like PyTorch and TensorFlow provide flags for deterministic behavior (e.g., torch.use_deterministic_algorithms(True)  [^1] https://pytorch.org/docs/stable/generated/torch.use_deterministic_algorithms.html). These ensure that certain operations avoid algorithms that could produce different results across runs. However, in practice, it can still produce non-deterministic results even with these flags .
Despite these efforts, achieving perfect reproducibility in LLM inference remains a challenge. The next subsection delves into a crucial factor that often leads to non-deterministic behavior even with greedy decoding and the aforementioned practices in place.
2.2 Numerical Precision and Rounding Errors
@@
Moreover, a key aspect of floating-point arithmetic that contributes to non-determinism is its non-associativity. This means that the order in which numbers are added can affect the final result due to accumulated rounding errors. Table 2 provides illustrative examples of how the order of summation can lead to different results in FP32 and BF16. This becomes particularly relevant in the parallel computations performed in GPUs during LLM inference.
Table 2: Two illustrative cases demonstrate how rounding error arising from the non-associativity of floating-point addition can affect numerical results. Example 1 reveals accumulation error at both precisions; Example 2 exhibits a discrepancy only in BF16, while FP32 delivers identical results, illustrating higher-precision numeric types are more tolerant of rounding errors.
@@
In the context of LLMs, even small numerical variations in the logit values can affect the final token selection when the top probabilities are close. Floating-point arithmetic in GPUs exhibits non-associativity, meaning $(a+b)+c\neq a+(b+c)$ due to finite precision and rounding errors. This property directly impacts the computation of attention scores and logits in the transformer architecture, where parallel operations across multiple threads can yield different results based on execution order. Formally, for a reduction operation $\oplus$ over a set of values $\{v_{1},v_{2},...,v_{n}\}$ (such as summing attention scores or computing softmax denominators), the result depends on the specific ordering of operations:
$\bigoplus_{i=1}^{n}v_{i}=v_{\pi(1)}\oplus v_{\pi(2)}\oplus...\oplus v_{\pi(n)},$
where $\pi$ is a permutation of indices $\{1,2,...,n\}$ determined by thread scheduling.
@@ 3.1 Experiment Setup
We conduct experiments on four recent LLMs, including two reasoning models: DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B and two non-reasoning models: Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct guo2025deepseek ; qwen2024 ; llama2024 , across five commonly used LLM evaluation benchmarks: AIME’24, MATH500, LiveCodeBench-Easy, LiveCodeBench-Medium, and LiveCodeBench-Hard aime2024 ; math500 ; livecodebench2024 . For reasoning models, we set the maximum output token length to 32,768 and for non-reasoning models, we set it to 2,048. We use vLLM vllm as the inference backend. In the random sampling experiments, we set temperature to 0.7 and top-p to 0.95. Our Evaluation implementation and prompt setting is adapted from SkyThought-evals sky_t1_2025 , more details can be found in Appendix C.
@@
Table 3: Std@Acc of greedy decoding across 12 different settings (GPU types, GPU counts, and batch sizes) under BF16, FP16, and FP32 Numerical Precisions. Reasoning models also exhibit larger variance compared to non-reasoning counterparts. More results can be found in Appendix E.
@@
Table 4: Standard deviation of output length of greedy decoding across 12 different settings (GPU types, GPU counts, and batch sizes) under BF16, FP16, and FP32 numerical precisions. The output length of reasoning models exhibit large variance. More results can be found in Appendix E.
@@ 3.2 Greedy Decoding $\neq$ Deterministic Output
Figure 5: Distribution of Div_Index for DeepSeek-R1-Distill-Qwen-7B on MATH500 under BF16, FP16, and FP32. Higher numerical precision leads to much fewer divergent examples and a shift of divergence onset to later token positions.
The impact of numerical precision on output stability is further evidenced by divergence patterns in greedy decoding. Figure 5 shows the distribution of divergence points (Div_Index) across precision formats for DeepSeek-R1-Distill-Qwen-7B on MATH500. As precision increases from BF16 to FP32, we observe both fewer divergent examples overall and a significant shift in when divergence occurs. With FP32, almost all examples produce identical outputs across configurations, resulting in only 2.2% of the samples diverged. In contrast, with BF16, divergence frequently occurs early in generation despite the deterministic setting, with over 90% of examples showing some form of divergence. When divergence happens in higher precision formats, it typically occurs much later in the sequence, limiting its impact on the final output and answer accuracy.
@@ 3.3 Random Sampling Has Reproducibility Issue, Too
One might argue that while greedy decoding is vulnerable to token-level instability from numerical precision, random sampling might be less sensitive due to its intrinsic stochasticity. However, our experiments reveal that numerical precision significantly affects the stability and reproducibility of sampling-based evaluations as well. When using random sampling with temperature $T>0$, researchers typically report the mean accuracy averaged over multiple runs (commonly referred to as Pass@1). We conduct Pass@1 evaluations across two benchmarks–AIME’24 and MATH500–using 16 and 64 independent sampling runs for AIME’24 and 4 runs for MATH500. Here we emphasize that in Table 5, the reported standard deviation is calculated based on Pass@1 performance across 6 system configurations per model (batch sizes and GPU counts), not across repeated runs of the same configuration. Thus, the reported variance largely reflects the impact from limited numerical precision, not the inherent variance of models. As shown in Table 5, we observe a clear trend: numerical precision introduces an additional source of variance beyond the intended randomness, and lower-precision formats such as BF16 tend to produce higher output variance. See Appendix H for the complete Pass@1 results that support these variance statistics.
@@ 3.4 Ablation: How Runtime Configurations Affect Reproducibility
Our analysis of runtime configurations reveals three key factors affecting token probability variations. First, in Figure 6 (a), configurations with 4 GPUs tend to exhibit higher probability variation than those with 2 GPUs across 3 tested batch sizes (particularly in BF16 precision), potentially due to increased parallel computation introducing more varied floating-point operation orderings and consequently different rounding errors. Second, Figure 6 (b) suggests that smaller batch sizes counter-intuitively produce higher variance in token probabilities because they may require more sequential processing steps that accumulate rounding errors, while larger batches benefit from parallel computation within optimized CUDA kernels that limit error accumulation. Third, GPU architecture matters: Figure 6 (c) shows A100s generally exhibit slightly higher probability variance than L40S under identical configurations, likely due to differences in hardware-level floating-point implementations and memory hierarchies. All these effects are most pronounced under BF16 precision with its limited mantissa bits making it especially susceptible to rounding effects. For more results of runtime configurations and tested tasks, please refer to the Appendix F.
@@ 4 Near-Perfect Deterministic Reproduction: LayerCast
Our experimental results strongly support this approach. When examining the standard deviation of accuracy across runs, Layer Cast achieves stability nearly identical to FP32, while BF16 shows much higher variability. As shown in Figure 8, the divergence index measurements further confirm that LayerCast produces consistent outputs across different batch sizes and GPU configurations, with divergence rates below 3.4%. From resource perspective, LayerCast offers substantial benefits over full FP32: memory usage is reduced by 34% (particularly important for KV cache in long-context scenarios). These improvements make LayerCast a practical solution for applications requiring both deterministic outputs and reasonable performance. For full result, please refer to Appendix G.
@@ 6 Conclusion
Appendix A Limitations
Despite the importance of our work in highlighting and addressing numerical precision’s effect on reproducibility, there are limitations that are worth noting. First, our experiments focus on a certain set of models and settings. Due to resource constraints, we do not conduct large-scale experiments on much larger models (>70B) or a wide range of GPU architectures or accelerators; these settings might exhibit different levels of numerical behaviors. Additionally, our study primarily addresses numerical precision issues in transformer-based LLMs and may not fully generalize to other architectures or modalities.
Appendix B Broader Impacts
This research has several societal impacts. First, by raising the discussion of reproducibility issues in LLM inference, our work promotes greater scientific rigor in AI research, calling for a more reliable comparison of models and techniques. This is particularly important as LLMs are increasingly deployed in critical areas like healthcare, education, and public services, where consistency and reproducibility are essential. On the other hand, our findings reveal that true reproducibility may require additional computational resources, potentially exacerbating the already significant environmental effects of LLM development. We hope our work will encouage the AI community to establity better standards of evaluation–ultimatelly leading to more trustworthy and reliable AI systems.
Appendix C Prompt Formats Used in Our Evaluation
@@
C.1 MATH Benchmarks: AIME’24 and MATH500
@@
C.2 Code Generation Benchmarks: LiveCodeBench-Easy, LiveCodeBench-Medium, LiveCodeBench-Hard
@@
Appendix D Supplementary Results on Accuracy Variance under BFloat16
@@
Appendix E Supplementary Results on Greedy Decoding
@@
Appendix F Supplementary Results on Ablation Study

### v2

# Understanding and Mitigating Numerical Sources of Nondeterminism in LLM Inference

@@ Abstract
Large Language Models (LLMs) are now integral across various domains and have demonstrated impressive performance. Progress, however, rests on the premise that benchmark scores are both accurate and reproducible. We demonstrate that the reproducibility of LLM performance is fragile: changing system configuration, such as evaluation batch size, GPU count, and GPU version, can introduce significant differences in the generated responses. This issue is especially pronounced in reasoning models, where minor rounding differences in early tokens can cascade into divergent chains of thought, ultimately affecting accuracy. For instance, under bfloat16 precision with greedy decoding, a reasoning model like DeepSeek-R1-Distill-Qwen-7B can exhibit up to 9% variation in accuracy and 9,000 tokens difference in response length due to differences in GPU count, type, and evaluation batch size. We trace the root cause of this variability to the non-associative nature of floating-point arithmetic under limited numerical precision. This work presents the first systematic investigation into how numerical precision affects reproducibility in LLM inference. Through carefully controlled experiments across various hardware, software, and precision settings, we quantify when and how model outputs diverge. Our analysis reveals that floating-point precision—while critical for reproducibility—is often neglected in evaluation practices. Inspired by this, we develop a lightweight inference pipeline, dubbed LayerCast, that stores weights in 16-bit precision but performs all computations in FP32, balancing memory efficiency with numerical stability. Code is available at https://github.com/nanomaoli/llm_reproducibility.
@@ 1 Introduction
We highlight the significance of this issue for several critical reasons. First, many researchers report benchmark performance based on a single inference run with a fixed random seed to reduce computational cost. However, as shown in Figure 1. This practice can lead to misleading conclusions about model performance. Second, even when performing random sampling with multiple independent runs, the averaged results can still vary significantly due to hardware and system-level nondeterminism. Moreover, when researchers report standard deviations without accounting for this numerical nondeterminism, they risk severely overestimating a model’s true uncertainty, since the reported variance reflects a mixture of intrinsic model uncertainty and variance introduced by finite numerical precision. When results cannot be exactly reproduced, it becomes difficult to distinguish whether improvements are from better methods or merely random variation.
@@
  • An optimized FP32 inference pipeline. We propose an optimized inference pipeline called LayerCast, which performs all computations in FP32 while retaining model weights in BF16 precision. This approach effectively balances memory efficiency and reproducibility. We release it as a patch to vLLM and can be used with just a few lines of code change.
2 Preliminary and Analysis
@@ 2.1 Current Practices on LLM Reproducibility
There are two widely adopted evaluation strategies. The first one is greedy decoding, where temperature is set to zero and model always select the token with highest probability. The second strategy employs random sampling with a non-zero temperature, and evaluates performance using the Pass@K metric chen2021evaluating . Below, we introduce the commonly adopted experimental setting for the reproducibility.
@@
Deterministic libraries and Random Seeds. Random seeds in LLM generation control the pseudo-random number generator that selects tokens when using non-greedy decoding. It is a standard practice to fix random seeds; however, this is not always sufficient for ensuring full determinism in greedy decoding. Even with a fixed seed, differences in computation order can alter the sequence of the pseudo-random number generator hochlehnert2025sober . Meanwhile, frameworks like PyTorch and TensorFlow provide flags for deterministic behavior (e.g., torch.use_deterministic_algorithms(True)  [^1] https://pytorch.org/docs/stable/generated/torch.use_deterministic_algorithms.html). These ensure that certain operations avoid algorithms that could produce different results across runs. However, in practice, it can still produce nondeterministic results even with these flags.
Despite these efforts, achieving perfect reproducibility in LLM inference remains a challenge. The next subsection delves into a crucial factor that often leads to nondeterministic behavior even with greedy decoding and the aforementioned practices in place.
2.2 Numerical Precision, Rounding Errors, and GPU Kernels
@@
Moreover, a key aspect of floating-point arithmetic that contributes to nondeterminism is its non-associativity. This means that the order in which numbers are added can affect the final result due to accumulated rounding errors. Table 2 provides illustrative examples of how the order of summation can lead to different results in FP32 and BF16. As the examples demonstrate, although rounding follows deterministic rules, non-associativity introduces nondeterminism, which is further amplified by the larger rounding error of BF16. This becomes particularly relevant in the parallel computations performed in GPUs during LLM inference.
Table 2: Two illustrative cases demonstrate how rounding error, together with the non‑associativity of floating‑point addition, can affect numerical results. Example 1 reveals accumulation error at both precisions; Example 2 exhibits a discrepancy only in BF16, while FP32 delivers identical results, illustrating higher‑precision numeric types are more tolerant of rounding errors.
@@
In the context of LLMs, even small numerical variations in the logit values can affect the final token selection when the top probabilities are close. Floating-point arithmetic in GPUs exhibits non-associativity, meaning $(a+b)+c\neq a+(b+c)$ due to finite precision and rounding errors. This property directly impacts the computation of attention scores and logits in the transformer architecture, where parallel operations across multiple threads can yield different results based on execution order. Formally, for a reduction operation $\oplus$ over a set of values $\{v_{1},v_{2},...,v_{n}\}$ (such as summing attention scores or computing softmax denominators), the result depends on the specific ordering of operations: $\bigoplus_{i=1}^{n}v_{i}=v_{\pi(1)}\oplus v_{\pi(2)}\oplus...\oplus v_{\pi(n)},$ where $\pi$ is a permutation of indices $\{1,2,...,n\}$ determined by thread scheduling.
GPU Kernels add numbers in different orders when changing system configuration. In serving systems, several factors can change the order in which floating-point operations are executed, including: (1) continuous batching yu2022orca , which dynamically modifies the set of requests within a batch; (2) different operator implementations, such as Split-K versus Non-Split-K MatMul NVIDIA_CUTLASS_EfficientGEMM ; (3) operator hyperparameters, like the block size in MatMul or FlashAttention; (4) collective operations in parallel settings, such as AllReduce; and (5) parallelization strategies, such as tensor parallelism (TP), which distribute computation across multiple GPUs. Together, all these factors can affect the determinism of LLM inference. To address this issue, Batch-invariant operations batchinv , including batch-invariant FlexAttention, RMSNorm, and MatMul kernels, are introduced. These kernels guarantee that inference results remain deterministic regardless of batch size variations. Specifically, this approach achieves determinism by parallelizing computation along the batch dimension, thereby decoupling batch size from the final outputs of individual requests. However, as the name “batch-invariant” suggests, the technique is currently limited to handling variations related only to the batch dimension, making it robust to continuous batching and other batch-size–related changes, but not to other forms of nondeterminism like changing the TP sizes or GPU types.
@@ 3.1 Experiment Setup
We conduct experiments on four recent LLMs, including two reasoning models: DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B and two non-reasoning models: Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct guo2025deepseek ; qwen2024 ; llama2024 , across five commonly used LLM evaluation benchmarks: AIME’24, MATH500, LiveCodeBench-Easy, LiveCodeBench-Medium, and LiveCodeBench-Hard aime2024 ; math500 ; livecodebench2024 . To verify the generalizability of our findings, we also conduct additional experiments on a larger model (Qwen3-32B) and a diverse reasoning benchmark (GPQA Diamond) covering graduate-level science questions, with results provided in the Appendix. For reasoning models, we set the maximum output token length to 32,768 and for non-reasoning models, we set it to 2,048. We primarily use vLLM vllm as the inference backend; to verify that our findings are not backend-specific, we also conduct verification experiments using HuggingFace Transformers (see Appendix for details). In the random sampling experiments, we set temperature to 0.7 and top-p to 0.95. Our Evaluation implementation and prompt setting is adapted from SkyThought-evals sky_t1_2025 , more details can be found in Appendix A.
@@
Table 3: Std@Acc of greedy decoding across 12 different settings (GPU types, GPU counts, and batch sizes) under BF16, FP16, and FP32 Numerical Precisions. Reasoning models also exhibit larger variance compared to non-reasoning counterparts. More results can be found in Appendix C.
@@
Table 4: Standard deviation of output length of greedy decoding across 12 different settings (GPU types, GPU counts, and batch sizes) under BF16, FP16, and FP32 numerical precisions. The output length of reasoning models exhibit large variance. More results can be found in Appendix C.
@@ 3.2 Greedy Decoding $\neq$ Deterministic Output
Figure 5: Distribution of Div_Index for DeepSeek-R1-Distill-Qwen-7B on MATH500 under BF16, FP16, and FP32. Higher numerical precisions lead to fewer divergent examples and a shift of divergence point to later token positions.
The impact of numerical precision on output stability is further evidenced by divergence patterns in greedy decoding. Figure 5 shows the distribution of divergence points (Div_Index) across precision formats for DeepSeek-R1-Distill-Qwen-7B on MATH500. As precision increases from BF16 to FP32, we observe both fewer divergent examples overall and a significant shift in when divergence occurs. With FP32, almost all examples produce identical outputs across configurations, resulting in only 2.2% of the samples diverging. In contrast, with BF16, divergence frequently occurs early in generation despite the deterministic setting, with over 90% of examples showing divergence. When divergence happens in higher precision formats, it typically occurs much later in the sequence, limiting its impact on the final output and answer accuracy.
@@ 3.3 Random Sampling Has Reproducibility Issue, Too
One might argue that while greedy decoding is vulnerable to token-level instability from numerical precision, random sampling might be less sensitive due to its intrinsic stochasticity. However, our experiments reveal that numerical precision significantly affects the stability and reproducibility of sampling-based evaluations as well. When using random sampling with temperature $T>0$, researchers typically report the mean accuracy averaged over multiple runs (commonly referred to as Pass@1). We conduct Pass@1 evaluations across two benchmarks–AIME’24 and MATH500–using 16 and 64 independent sampling runs for AIME’24 and 4 runs for MATH500. Here we emphasize that in Table 5, the reported standard deviation is calculated based on Pass@1 performance across 6 system configurations per model (batch sizes and GPU counts), not across repeated runs of the same configuration. Thus, the reported variance largely reflects the impact from limited numerical precision, not the inherent variance of models. As shown in Table 5, we observe a clear trend: numerical precision introduces an additional source of variance beyond the intended randomness, and lower-precision formats such as BF16 tend to produce higher output variance. See Appendix E for the complete Pass@1 results that support these variance statistics.
@@ 3.4 Ablation: How Runtime Configurations Affect Reproducibility
Our analysis of runtime configurations reveals three key factors affecting token probability variations. First, in Figure 6 (a), configurations with 4 GPUs tend to exhibit higher probability variation than those with 2 GPUs across 3 tested batch sizes (particularly in BF16 precision), potentially due to increased parallel computation introducing more varied floating-point operation orderings and consequently different rounding errors. Second, Figure 6 (b) suggests that smaller batch sizes counter-intuitively produce higher variance in token probabilities because they may require more sequential processing steps that accumulate rounding errors, while larger batches benefit from parallel computation within optimized CUDA kernels that limit error accumulation. Third, GPU architecture matters: Figure 6 (c) shows A100s generally exhibit slightly higher probability variance than L40S under identical configurations, likely due to differences in hardware-level floating-point implementations and memory hierarchies. All these effects are most pronounced under BF16 precision with its limited mantissa bits making it especially susceptible to rounding effects. For more results of runtime configurations and tested tasks, please refer to the Appendix D.
@@ 4 Near-Perfect Deterministic Reproduction: LayerCast
Our experimental results strongly support this approach. When examining the standard deviation of accuracy across runs, Layer Cast achieves stability nearly identical to FP32, while BF16 shows much higher variability. As shown in Figure 8, the divergence index measurements further confirm that LayerCast produces consistent outputs across different batch sizes and GPU configurations, with divergence rates below 3.4%. From resource perspective, LayerCast offers substantial benefits over full FP32: memory usage is reduced by 34% (particularly important for KV cache in long-context scenarios). These improvements make LayerCast a practical solution for applications requiring both deterministic outputs and reasonable performance. For full result, please refer to Appendix F.
@@ 6 Conclusion
Appendix A Prompt Formats Used in Our Evaluation
@@
A.1 MATH Benchmarks: AIME’24 and MATH500
@@
A.2 Code Generation Benchmarks: LiveCodeBench-Easy, LiveCodeBench-Medium, LiveCodeBench-Hard
@@
Appendix B Supplementary Results on Accuracy Variance under BFloat16
@@
Appendix C Supplementary Results on Greedy Decoding
@@
C.1 Additional Results on Larger Models and Diverse Tasks
To further validate our findings on the impact of numerical precision on reproducibility, we conducted additional experiments on a larger model (Qwen3-32B) and a diverse reasoning task (GPQA Diamond). These results demonstrate that the reproducibility challenges we identified are not limited to 7-8B models or math/coding tasks.
Larger Model Evaluation (Qwen3-32B): To demonstrate the impact of numerical precision on reproducibility for larger models, we benchmarked Qwen3-32B on the AIME’24 dataset under different runtime configurations covering 2 GPU counts (2 and 4), 2 GPU versions (A100 and L40S), and 3 batch sizes (8, 16, 32). Due to memory constraints, Qwen3-32B cannot be loaded into 2 L40S GPUs with FP32 precision, resulting in 9 feasible configurations for FP32 out of the 12 total configurations. The results in Table 10 confirm that larger models exhibit the same reproducibility challenges we identified in 7-8B models. Across all configurations, BF16 shows significant variance with Std@Acc of 5.02%, while FP32 maintains much better consistency with Std@Acc of only 1.11%.
Table 10: Accuracy of Qwen3-32B on AIME’24 under different runtime configurations (GPU types, GPU counts, and batch sizes) and numerical precisions. OOM indicates out-of-memory errors.
4 GPUs 2 GPUs
Precision A100-BS8 A100-BS16 A100-BS32 A100-BS8 A100-BS16 A100-BS32
BF16 80.00% 83.33% 83.33% 76.67% 83.33% 73.33%
FP32 76.67% 76.67% 76.67% 76.67% 80.00% 76.67%
4 GPUs 2 GPUs
Precision L40S-BS8 L40S-BS16 L40S-BS32 L40S-BS8 L40S-BS16 L40S-BS32
BF16 73.33% 83.33% 70.00% 83.33% 76.67% 73.33%
FP32 76.67% 76.67% 76.67% OOM OOM OOM
Diverse Task Evaluation (GPQA Diamond): To demonstrate that reproducibility issues extend beyond math and coding tasks, we evaluated DeepSeek-R1-Distill-Llama-8B on the GPQA Diamond benchmark, which focuses on graduate-level science questions. Table 11 shows the results across 12 different runtime configurations. The results reveal that BF16 exhibits Std@Acc of 3.03% compared to 1.96% for FP32, confirming that numerical precision affects reproducibility across diverse reasoning domains.
Table 11: Accuracy of DeepSeek-R1-Distill-Llama-8B on GPQA Diamond under different runtime configurations (GPU types, GPU counts, and batch sizes) and numerical precisions.
4 GPUs 2 GPUs
Precision A100-BS8 A100-BS16 A100-BS32 A100-BS8 A100-BS16 A100-BS32
BF16 33.84% 38.38% 41.41% 40.91% 41.92% 41.41%
FP32 40.91% 45.45% 43.43% 40.91% 43.43% 40.91%
4 GPUs 2 GPUs
Precision L40S-BS8 L40S-BS16 L40S-BS32 L40S-BS8 L40S-BS16 L40S-BS32
BF16 43.43% 38.38% 44.95% 38.89% 43.94% 41.92%
FP32 38.89% 42.42% 44.95% 42.93% 41.41% 44.44%
These additional experiments on a larger model and a diverse reasoning task further validate our core findings: numerical precision critically impacts reproducibility across model scales and task domains, with BF16 consistently showing higher variance compared to FP32.
C.2 Verification with HuggingFace Transformers Backend
To verify that the reproducibility issues we observed are not specific to the vLLM inference backend, we conducted additional experiments using the standard HuggingFace Transformers implementation without vLLM or PagedAttention. The experimental setup is identical to our main paper: 2 GPU counts (2 and 4), 2 GPU versions (A100 and L40S), and 3 batch sizes (8, 16, 32), covering 12 settings for each precision.
Table 12 shows the results for two models on AIME’24. The same pattern emerges: FP32 inference is fully reproducible with 0% Std@Acc for both models, while BF16 exhibits Std@Acc of 2.23% for Qwen2.5-7B-Instruct and 1.51% for Llama-3.1-8B-Instruct, similar variance to what we reported with vLLM. These results demonstrate that the reproducibility phenomenon is fundamental to GPU hardware and floating-point arithmetic, and is observable across different inference backends.
Table 12: Std@Acc using HuggingFace Transformers backend (without vLLM) on AIME’24 across 12 different settings, confirming that reproducibility issues are not backend-specific.
Model / Dataset Std@Acc BF16 Std@Acc FP32
Qwen2.5-7B-Instruct / AIME’24 2.23% 0
Llama-3.1-8B-Instruct / AIME’24 1.51% 0
Appendix D Supplementary Results on Ablation Study